The amount of time spent by North American adults watching television per day is normally distributed with a mean of 6 hours and a standard deviation of 1.5 hours. (a) What proportion of the population watches television for more than 7 hours per day? (b) What is the probability that the average number of hours spent watching television by a random sample of five adults is more than 7 hours? (c) What is the probability that in a random sample of five adults all watch television for more than 7 hours per day?


1
Expert's answer
2022-04-20T00:20:00-0400

Let "X=" the amount of time spent by North American adults watching television per day: "X\sim N(\mu, \sigma^2)."

Then "Z=\dfrac{X-\mu}{\sigma}\sim N(0, 1)"

a) Given "\mu=6\ hours, \sigma=1.5 \ hours"



"P(X>7)=1-P(X\leq7)""=1-P(Z\leq\dfrac{7-6}{1.5})\approx 1-P(Z<0.666667)""\approx1-0.747507\approx0.252493"

b) Given "\mu=6\ hours, \sigma=1.5 \ hours, n=5"



"X\sim N(\mu, \sigma^2/n), Z=\dfrac{X-\mu}{\sigma/\sqrt{n}}\sim N(0, 1)""P(X>7)=1-P(X\leq7)""=1-P(Z\leq\dfrac{7-6}{1.5/\sqrt{5}})\approx 1-P(Z<1.490712)""\approx1-0.931981\approx0.068019"

b) Given "\mu=6\ hours, \sigma=1.5 \ hours, n=30"



"X\sim N(\mu, \sigma^2/n), Z=\dfrac{X-\mu}{\sigma/\sqrt{n}}\sim N(0, 1)""P(X>7)=1-P(X\leq7)""=1-P(Z\leq\dfrac{7-6}{1.5/\sqrt{5}})\approx 1-P(Z<1.490712)""\approx1-0.931981\approx0.068019"

c) The likelihood that five out of five adults watch TV more than 7 hours a day is the first probability raised to the fifth power



"(0.252493)^5 \approx0.001026"